# Oscillatrons: neural units with time-dependent multifunctionality

### v2

@@ Abstract
Several branches of computing use a system’s physical dynamics to do computation. We show that the dynamics of an underdamped harmonic oscillator can perform multifunctional computation, solving distinct problems at distinct times within a dynamical trajectory. Oscillator computing usually focuses on the oscillator’s phase as the information-carrying component. Here we focus on the time-resolved amplitude of an oscillator whose inputs influence its frequency, which has a natural parallel as the activity of a time-dependent neural unit. We call this unit an oscillatron. The activity of an oscillatron at fixed time is a nonmonotonic function of the input, and so it can solve nonlinearly-separable problems such as XOR. The activity of the oscillatron at fixed input is a nonmonotonic function of time, and so it is multifunctional in a temporal sense, able to carry out distinct nonlinear computations at distinct times within the same dynamical trajectory. Time-resolved computing of this nature can be done in or out of equilibrium, with the natural time evolution of the system giving us multiple computations for the price of one.
@@ I Introduction
Here we show that the explicit time dependence of a physical dynamics permits multifunctional computation, with a single device able to perform multiple distinct calculations in the course of a dynamical trajectory. To illustrate this idea we consider the time-dependent dynamics of a continuous-valued underdamped oscillator. Oscillators can be realized experimentally in many ways Von Neumann (1957); Csaba and Porod (2020); Ciliberto (2017), including by mechanical cantilevers Dago et al. (2021, 2022) and electrical circuits Wang and Roychowdhury (2019); Chou et al. (2019); Melanson et al. (2023). Computing with oscillators is a concept that dates back to the 1950s Von Neumann (1957); Goto (1959). Most examples of oscillator-based computing focus on the phase of the oscillator as a means of carrying information, and aim to find oscillatory ground states in networks of interacting oscillators Csaba and Porod (2020); Bonnin et al. (2022).
@@
Here we consider the time-resolved amplitude of an oscillator whose inputs influence its frequency, which has a natural parallel as the activity of a time-dependent neural unit. The motivation for this choice is twofold. First, because the activity of the unit at fixed time is a nonmonotonic function of the input, a single unit can solve nonlinearly-separable problems such as XOR, making it more expressive than standard artificial neurons. Second, because the activity of the unit at fixed input is a nonmonotonic function of time, the unit is multifunctional in a temporal sense, able to carry out different nonlinear computations at different times within the same dynamical trajectory.
Neural units that are a nonmonotonic function of their input are more expressive than standard artificial neurons. For example, logic operations such as XOR cannot rendered linearly separable by a standard perceptron unit, which applies a monotonic activation function to a linear function of its input data Rumelhart et al. (1986), but can be solved by oscillator units Gidon et al. (2020); Noel et al. (2021). Consider two binary variables, $I_{1,2}=0,1$: the oscillatory function $f(I_{1},I_{2})=\sin(\pi(I_{1}+I_{2})/2)$ makes the XOR problem linearly separable, because $f(0,0)=f(1,1)=0$ and $f(0,1)=f(1,0)=1$. Some neurons in the brain operate in an oscillatory way Gidon et al. (2020), an observation that has motivated other authors to show that neural networks built from oscillator units are highly expressive Effenberger et al. (2022).
Moreover, temporal oscillations permit a single dynamical neural unit to be multifunctional, performing multiple computations in the course of a single dynamical trajectory. We show that a single oscillator neuron can act as all of the elementary logic gates, depending on the time at which we measure its output, and can be trained by gradient descent to do distinct classification tasks at distinct times. A device built from such units could perform multiple computations in a single dynamical trajectory, requiring only a single set of parameters to do multiple tasks.
@@
A neural unit whose time-resolved activity is the information-carrying element and which oscillates with a frequency influenced by its inputs we call an oscillatron, by comparison with perceptron, whose activity depends in a monotonic way on its inputs. Related concepts include the parametron, an oscillating neural unit whose phase is the information-carrying element Goto (1959), and the Oscillatron-1, a neural-network architecture whose pattern generators oscillate in synchrony with its feature detectors Atkins (1991).
Figure 1: (a) The time-resolved dynamics (2) of the oscillatory neural unit (1), or oscillatron, evaluated for the two distinct values of a binary input $I=0,1$. Depending on the observation time $t$, the unit can perform any of the elementary one-bit operations. (b) The time-resolved dynamics of the same neuron, now evaluated for the three distinct values $I=I_{1}+I_{2}=0,1,2$ of the binary inputs $(I_{1},I_{2})$. Depending on the observation time $t$, the unit can function as any of the elementary logic gates. Neuron parameters: $\omega_{0}=1,\gamma=1/10$.
II Oscillatrons are multifunctional in a time-dependent way
@@
In more detail, consider a continuous degree of freedom $S(t)$ of unit mass that evolves according to the underdamped Langevin dynamics $\ddot{S}=-\gamma\dot{S}-\partial_{S}U+\eta(t)$. Here $\gamma$ is a friction coefficient, $U$ is a potential, and $\eta$ is a noise term. We will consider the harmonic potential $U=\frac{1}{2}(\omega_{0}^{2}+I)S^{2}$, where $\omega_{0}^{2}$ is the fundamental frequency of the oscillator and $I$ is the input signal (the input signal could in general be time-dependent, but here we consider a constant input). The input therefore influences the spring constant of the harmonic potential. In general, thermal noise provides an important mechanism for driving dynamical evolution and allowing probabilistic computation Aifer et al. (2023). Here, for simplicity, we assume the low-noise limit, and so $S(t)$ evolves according to the equation
@@
We assume initial conditions $S(0)=1$ and $\dot{S}(0)=0$. Then $S$ satisfies
@@
is a function of the intrinsic parameters $\omega_{0}$ and $\gamma$ and the input $I$. We assume the underdamped regime, where $\Omega^{2}>0$, in which case $S$ oscillates with frequency $\Omega$. At fixed time the output of the unit is a nonmonotonic function of $I$, and at fixed $I$ the output of the unit is a nonmonotonic function of $t$. These two properties give this neural unit the ability to solve nonlinearly separable problems, and to do so in a temporally multifunctional way. $S$ is a damped harmonic oscillator whose input $I$ influences its frequency of oscillation; in a computing context we call $S$ an oscillatron.
@@
An oscillatron can perform computations in a multifunctional way. Consider a binary input $I=0,1$. In Fig. 1(a) we show the neuron output (2) as a function of time $t$ for the two possible values of $I$. With a horizontal decision boundary placed at zero (and values above and below the boundary considered to be states 1 and 0, respectively), the neuron can perform any of the elementary one-bit operations, depending on observation time: identity ($I\to I$) at $t_{1}$; erase to one ($I\to 1$) at $t_{2}$; NOT or invert ($I\to\bar{I}$) at $t_{3}$; and erase to zero ($I\to 0$) at $t_{4}$. The same operations are also performed at other times in the unit’s trajectory.

### v3

# Neural units with time-dependent functionality

@@ Abstract
We show that the time-resolved dynamics of an underdamped harmonic oscillator can be used to do multifunctional computation, performing distinct computations at distinct times within a single dynamical trajectory. We consider the amplitude of an oscillator whose inputs influence its frequency. The activity of the oscillator at fixed time is a nonmonotonic function of its inputs, and so it can solve problems such as XOR that are not linearly separable. The activity of the oscillator at fixed input is a nonmonotonic function of time, and so it is multifunctional in a temporal sense, able to carry out distinct nonlinear computations at distinct times within the same dynamical trajectory. We show that a single oscillator, observed at different times, can act as all of the elementary logic gates and can perform binary addition, the latter usually implemented in hardware using 5 logic gates. We show that a set of $n$ oscillators, observed at different times, can perform an arbitrary number of analog-to-$n$-bit digital conversions. We also show that oscillators can be trained by gradient descent to perform distinct classification tasks at distinct times. Computing with time-dependent functionality can be done in or out of equilibrium, and suggests a way of reducing the number of parameters or devices required to do nonlinear computations.
@@ I Introduction
Here we show that the explicit time dependence of a physical dynamics can be used to do multifunctional computation, with a single device able to perform multiple distinct calculations in the course of a dynamical trajectory. We consider the dynamics of a continuous-valued underdamped oscillator. Oscillators can be realized in hardware in many ways Neumann (1957); Csaba and Porod (2020); Ciliberto (2017), including by springs French (1971), cantilevers Dago et al. (2021, 2022), and electrical circuits Harris and Harris (2015); Mano (2017); Wang and Roychowdhury (2019); Chou et al. (2019); Melanson et al. (2023). Computing with oscillators is a concept that dates back to the 1950s Neumann (1957); Goto (1959). Most examples of oscillator-based computing focus on the phase of the oscillator as a means of carrying information, and aim to find oscillatory ground states in networks of interacting oscillators Csaba and Porod (2020); Bonnin et al. (2022).
Figure 1: An illustration of the computational power of neural units with nonmonotonic activations, using binary inputs $I_{1,2}=0,1$. (a) A standard perceptron unit, whose output is a monotonic function of a linear sum of its inputs, can solve AND. (b) However, it cannot solve XOR, which is not linearly separable: neither red line (nor any other straight line) separates the two classes. (c) A multilayer perceptron is required to solve XOR: its hidden layers transform the input data $(I_{1},I_{2})\to(\tilde{I}_{1},\tilde{I}_{2})$ until a linear decision boundary suffices. (d) Alternatively, XOR can be solved by a neural unit whose output is a nonmonotonic function of its inputs Vodenicarevic et al. (2017); Gidon et al. (2020); Noel et al. (2021), here $f(I_{1},I_{2})=\sin(\pi(I_{1}+I_{2})/2)$. In this paper we build on this observation to show that the dynamics of an underdamped oscillator, whose output is a nonmonotonic function of its inputs and the elapsed time, can perform distinct nonlinear computations at distinct times.
@@
Here we focus on the time-resolved amplitude of an oscillator whose inputs influence its frequency. The motivation for this choice is twofold. First, because the activity of the oscillator at fixed time is a nonmonotonic function of the input, a single oscillator can solve problems such as XOR that are not linearly separable, making it more expressive than standard artificial neurons. Second, because the activity of the oscillator at fixed input is a nonmonotonic function of time, it is multifunctional in a temporal sense, able to carry out different nonlinear computations at different times within the same dynamical trajectory.
Units that are a nonmonotonic function of their inputs are more expressive than standard artificial neurons. For example, the standard perceptron unit, which applies a monotonic activation function to a linear function of its inputs, can solve problems that are linearly separable, such as the AND operation Rumelhart et al. (1986). Consider two binary variables, $I_{1,2}=0,1$. The AND operation returns 1 if both inputs are 1, and returns 0 otherwise. As shown in Fig. 1(a), a single decision boundary that is a linear function of the two inputs can separate the two output classes, and so a perceptron can solve AND.
The same is not true of the XOR operation. XOR returns 0 if $I_{1}$ and $I_{2}$ are equal, and returns 1 otherwise. As shown graphically in Fig. 1(b), there is no linear function of the input data for which a monotonic activation function can separate the classes 0 and 1, and so a single perceptron cannot solve XOR. The XOR problem demonstrated the limitations of single-layer perceptrons, leading to the development of multi-layer neural networks and the field of deep learning LeCun et al. (2015); Schmidhuber (2015). As sketched in Fig. 1(c), a multilayer perceptron transforms the input data so that a final layer can implement a linear decision boundary that solves the problem.
Alternatively, XOR can be solved by a single oscillatory unit Vodenicarevic et al. (2017); Gidon et al. (2020); Noel et al. (2021). The oscillatory function $f(I_{1},I_{2})=\sin(\pi(I_{1}+I_{2})/2)$, shown as a color map in Fig. 1(d), satisfies $f(0,0)=f(1,1)=0$ and $f(0,1)=f(1,0)=1$ and so solves XOR.
Thus a neural unit that is a non-monotonic function of its output is in this sense more powerful than a monotonic activation function. Some neurons in the brain display an output that is a non-monotonic function of their input, enabling them to solve XOR Gidon et al. (2020). Such observations have motivated other authors to show that neural networks built from oscillator units are highly expressive: in Ref. Effenberger et al. (2022), a recurrent neural network built from damped harmonic oscillators outperformed similar networks built from non-oscillating units, in both learning speed and task performance. While biological neurons and harmonic oscillators are different in detail – for instance, biological neurons are stochastic, exhibit plasticity, and undergo various biochemical processes that include growth and aging – an oscillator model can serve as a simplified abstraction of the non-monotonic behavior observed in biological neurons.
Here we build on these observations by showing that the dynamics of an underdamped oscillator, whose output is a nonmonotonic function of its inputs and the elapsed time, can perform distinct nonlinear computations at distinct times. In Section II we show that a single oscillator unit can act as all of the elementary logic gates, depending on the time at which we measure its output, and can do binary arithmetic conventionally done using multiple logic gates. We also show that $n$ oscillators used in parallel, observed at different times, can perform an arbitrary number of analog-to-$n$-bit digital conversions. In Section III we show that oscillators can be trained, by gradient-based adjustment of their input weights, to recognize distinct image types at distinct times. We conclude in Section IV.
@@
A neural unit whose time-resolved activity is the information-carrying element and which oscillates with a frequency influenced by its input could be called an oscillatron  [^1] The Oscillatron-1 is neural network whose pattern generators oscillate in synchrony with its feature detectors Atkins (1991)., by comparison with perceptron, the standard monotonic neural unit, and parametron, an oscillating neural unit whose phase is the information-carrying element Goto (1959). Here we will use the terms oscillator, oscillatron, and neuron interchangeably. A device built from such units could perform multiple computations within a single dynamical trajectory, requiring only a single set of parameters to do multiple tasks.
II Dynamical oscillators possess time-dependent functionality
Consider a continuous degree of freedom $S(t)$ of unit mass that evolves according to the underdamped Langevin dynamics Van Kampen (1992)
$\ddot{S}=-\gamma\dot{S}-\partial_{S}U(S)+\eta(t).$ (1)
@@ II Dynamical oscillators possess time-dependent functionality
Here $\gamma$ is a friction coefficient and $U$ is a potential. We will consider the harmonic potential $U(S)=\frac{1}{2}(\omega_{0}^{2}+I)S^{2}$, where $\omega_{0}$ is an intrinsic frequency and $I$ is an input signal. The input signal could in general be time-dependent, but here we consider a constant input (for instance, constant connections between RLC oscillators can be engineered in circuit boards Aifer et al. (2023)). The input therefore influences the spring constant of the harmonic potential. The term $\eta(t)$ in (1) describes thermal noise, which in general provides an important mechanism for driving dynamical evolution and allowing probabilistic computation Aifer et al. (2023). Here we assume the low-noise limit, and so $S(t)$ evolves according to the equation
@@
We impose the initial conditions $S(0)=1$ and $\dot{S}(0)=0$, and so
@@
is a function of the input $I$ and the oscillator’s intrinsic parameters $\omega_{0}$ and $\gamma$
Figure 2: (a) The time-resolved dynamics (3) of the oscillatory neural unit (2), or oscillatron, evaluated for the two distinct values of a binary input $I=0,1$. Depending on the observation time $t$, the unit can perform any of the elementary one-bit operations. (b) The time-resolved dynamics of the same unit, now evaluated for the three distinct values $I=I_{1}+I_{2}=0,1,2$ of the binary inputs $(I_{1},I_{2})$. Depending on the observation time $t$, the unit can function as any of the elementary logic gates. (c) A binary full adder realized by the unit (2) with input $I=A+B+C_{\rm in}$. We show time traces $S_{I}(t)$, Eq. (3), for the four possible values of $I$. The sum $\Sigma$ and carry-out $C_{\rm out}$ are given by the output of the unit at times $t_{1}$ and $t_{2}$, respectively. Oscillator parameters: $\omega_{0}=1,\gamma=1/10$.
@@
We consider the underdamped regime, where $\Omega^{2}>0$, in which case $S$ oscillates with frequency $\Omega$. $S$ is a damped harmonic oscillator whose input $I$ influences its frequency of oscillation, which in computing terms could be called an oscillatron. At fixed time $t$ the output of the unit is a nonmonotonic function of $I$, and at fixed $I$ the output of the unit is a nonmonotonic function of $t$. These two properties give this neural unit the ability to solve problems that are not linearly separable, and to solve distinct problems of this nature at distinct times.
II.1 Logic gates
@@
As a first example, consider the binary input $I=0,1$. In Fig. 2(a) we show the oscillator output (3) as a function of time $t$ for the two possible values of $I$. With a horizontal decision boundary placed at zero (and values above and below the boundary considered to be states 1 and 0, respectively), the neuron can perform any of the elementary one-bit operations, depending on observation time: identity ($I\to I$) at $t_{1}$; set (or erase) to one ($I\to 1$) at $t_{2}$; NOT or invert ($I\to\bar{I}$) at $t_{3}$; and set to zero ($I\to 0$) at $t_{4}$. The same operations are also performed at other times in the unit’s trajectory.
Conventional physical models of processes such as bit erasure consider the manipulation of a degree of freedom in an external potential, with work expended Landauer (1961); Sagawa (2012); Zulkowski and DeWeese (2014); Proesmans et al. (2020); Dago and Bellon (2022). Here, by contrast, energy must be input in order to set the initial state of the unit and impose and maintain the input to the neuron (which sets the potential spring constant), but after setting the initial condition no work is performed: computations are done by the system’s natural dynamics. Moreover, the same unit can perform different computations when observed at different times.
Now consider two binary input degrees of freedom $I_{1,2}=0,1$, and construct the input $I=I_{1}+I_{2}$. (This can be done in hardware using an operational amplifier in summing configuration; the resulting signal sets the spring constant of the oscillator.) $I$ can take three values, 0, 1, or 2. In Fig. 2(b) we show the oscillator output (3) as a function of time $t$ for the three possible values of $I$. With the horizontal decision boundary placed as shown, the neuron can function as any of the elementary logic gates, depending on observation time. If we observe the unit at time $t_{1}$ it functions as an AND gate; at time $t_{2}$, an OR gate; and at time $t_{3}$, an XOR gate. At different times the same unit can also function as the inverted versions of these gates, NAND, XNOR, and NOR, the latter shown at time $t_{4}$.
Here we have chosen decision boundaries and observation times by inspection, given a choice of oscillator parameters. In Section III we show how to train oscillator parameters by gradient descent, given a choice of decision boundary and observation times.
Note that the equilibrium state of the unit is $S(t\to\infty)=0$, regardless of the value of $I$, and so in equilibrium the unit has no computational ability. In general, we can do time-resolved computation in or out of equilibrium, but doing it out of equilibrium means that we do not need to wait for equilibrium to be attained or verify that it has been attained.
II.2 Binary arithmetic
Binary arithmetic on a CPU is done by multiple logic gates Harris and Harris (2015); Mano (2017). For instance, a binary half adder turns binary inputs $A=0,1$ and $B=0,1$ into a sum $\Sigma=A\oplus B$, where $\oplus$ denotes the XOR operation, and a carry $C=A\cdot B$, where $\cdot$ denotes the AND operation. Consequently, the half adder is realized in hardware by two gates, XOR and AND Harris and Harris (2015); Mano (2017). Fig. 2(b) shows that one oscillator can represent both XOR and AND, at different observation times, indicating that it can function as a half adder: the sum and the carry are the output of the unit at times $t_{3}$ and $t_{1}$, respectively. To produce a steady binary signal the output of the unit must be fed to a comparator (to produce a binary signal from an analog one) and a clamp (to produce a time-invariant signal), but the nonlinear computation can be done by a single oscillator.
This idea can be generalized to more complex operations. A binary full adder turns binary inputs $A$ and $B$ and a carry-in $C_{\rm in}$ into a sum $\Sigma=A\oplus B\oplus C_{\rm in}$ and a carry-out $C_{\rm out}=(A\cdot B)+(B\cdot C_{\rm in})+(C_{\rm in}\cdot A)$, where $+$ denotes the OR operation Harris and Harris (2015); Mano (2017). The resulting truth table is
A B $C_{\rm in}$ $C_{\rm out}$ $\Sigma$
0 0 0 0 0
0 0 1 0 1
0 1 0 0 1
0 1 1 1 0
1 0 0 0 1
1 0 1 1 0
1 1 0 1 0
1 1 1 1 1
Table 1: Binary full adder: carry-out and sum for the inputs $A$, $B$, and $C_{\rm in}$.
The full adder is usually realized in hardware by 5 gates, two half adders and an OR gate Harris and Harris (2015); Mano (2017).
In terms of the four possible values of the combination $I=A+B+C_{\rm in}$, the truth table above can be written
$I$ $C_{\rm out}$ $\Sigma$
0 0 0
1 0 1
2 1 0
3 1 1
Table 2: Binary full adder: carry-out and sum for the combination $I=A+B+C_{\rm in}$.
This truth table can be represented by a single oscillator. In Fig. 2(c) we show the output $S_{I}(t)$, Eq. (3), of the unit (2), for the four possible values of its input $I=A+B+C_{\rm in}$. The output of the unit at time $t_{1}$ is the required sum, i.e. $\Sigma=S_{I}(t_{1})$, while the output of the unit at time $t_{2}$ is the required carry-out, i.e. $C_{\rm out}=S_{I}(t_{2})$. Thus one oscillatory unit can perform a computation usually done with 5 logic gates. Again, the output of the oscillator is analog and time-varying, and so must be fed to a comparator and latch in order to provide a steady binary signal, but the computation itself can be done by one device.
Harmonic oscillators can therefore do multi-bit binary addition: an $n$-bit adder can be built by chaining $n$ full adder devices in series, with the carry-out of each full adder sent to the carry-in of the next full adder. In this case a comparator and latch can function as an encoder, converting the output of one oscillator into a steady digital signal. An operational amplifier in summing configuration can then be used as an encoder, combining that signal with two additional inputs to set the spring constant of the next oscillator in the chain.
II.3 Multifunctional analog-to-digital conversion
In this section we consider a form of analog-to-digital conversion (ADC), with $n$ oscillators used to convert an analog input $A$ into an $n$-bit binary number. The same set of oscillators, observed at $m$ different times, can convert $m$ analog inputs into binary outputs. This section differs from the rest of the paper in that we consider the output of a set of oscillators whose fundamental frequencies are fixed and do not depend on the input, at observation times that are a function of the input.
Consider an analog number $A\in[0,A_{\rm max}]$, and form the input
$I(A)=\frac{A}{A_{\rm max}}(2^{n}-1).$ (5)
Here $n$ is the number of binary digits we wish to use to represent $A$. (A linear scaling can be done in hardware using a scaling amplifier or a shift register.) Table 3 shows the binary representation of the input $I$ sampled at 16 integer values for the case $n=4$, large enough to identify the following pattern: as $I$ increases, the value of each bit $B_{i}$ oscillates between 0 and 1 with period $2^{i}$.
$I$ $B_{4}$ $B_{3}$ $B_{2}$ $B_{1}$
0 0 0 0 0
1 0 0 0 1
2 0 0 1 0
3 0 0 1 1
4 0 1 0 0
5 0 1 0 1
6 0 1 1 0
7 0 1 1 1
8 1 0 0 0
9 1 0 0 1
10 1 0 1 0
11 1 0 1 1
12 1 1 0 0
13 1 1 0 1
14 1 1 1 0
15 1 1 1 1
Table 3: Binary digits $B_{i}=0,1$ representing an analog input $I$ sampled at integer values. Moving from top to bottom, the period of digit $B_{i}$ is $2^{i}$ (the red and blue each indicate a half period), and so can be represented by an oscillator with frequency $\omega_{i}=2\pi/2^{i}$.
This pattern indicates that the required bit value can be represented by an oscillator with the appropriate frequency, as shown in Fig. 3. The dashed vertical black line shows the binary representation of an input $1514\leq I<$ 15 to be 1110.
Figure 3: A graphical representation of Table 3, showing that the $i^{\rm th}$ binary digit $B_{i}$ of the analog input $I$ can be represented by an oscillator of period $2^{i}$, as $B_{i}=\Theta\left[S_{i}(t_{I})\right]$.
A set of oscillators of the type (2) can represent the required set of outputs. Denoting an oscillator’s fundamental frequency as $\omega_{i}$, and setting $\gamma=0$ for simplicity, we impose the initial conditions $S(0)=0$ and $\dot{S}(0)=-1$. Then (3) is replaced by
$S_{i}(t)=-\sin(\omega_{i}t).$ (6)
Consider $n$ such oscillators $i=1,2,\dots,n$, with fundamental frequencies $\omega_{i}=2^{1-i}\pi\,\omega_{0}$. With time measured in units of $\omega_{0}^{-1}$, the period of oscillator $i$ is $2^{i}$, and so, by the arguments above, the output of oscillator $i$ at time $t_{I}=I(A)$, interpreted as 0 or 1 if less than or greater than zero, gives the $i^{\rm th}$ binary digit of the scaled analog input $I(A)$. That is,
$B_{i}=\Theta\left[S_{i}(t_{I})\right],$ (7)
where $\Theta(x)$ is the Heaviside function, equal to 1 if $x>0$ and equal to zero otherwise. To this prescription we add the following caveat: if $I$ is an integer, we must observe the output of the oscillator at time $t_{I}=I(A)+\epsilon$, where $10<\epsilon<$ 1: the vertical black line in Fig. 3 indicates an observation time suitable for providing the binary representation of 14, namely 1110.
It is straightforward to verify that (7), with $i=0,1,2,3$, reproduces the entries $B_{i}$ of Table 3 when evaluated at integer $I$. But the formula also applies more generally, for non-integer $I$ and arbitrary bit number $n$.
Each bit is calculated independently, and does not require digits to be carried between units. The calculation requires $n$ oscillators and $n$ comparators, and requires an observation time of up to $12^{n}-$ 1, in units of $\omega_{0}^{-1}$  [^2] Note that oscillators made from RLC circuits can operate at frequencies $\omega_{0}$ of order GHz, comparable to the rate of a standard clock cycle on a CPU.. If we consider only a single analog-to-digital conversion, then this approach is less efficient than a successive approximation register, which requires one comparator and operates in a time that scales linearly with the number of bits $n$ (flash analog-to-digital conversion operates in constant time, but requires $12^{n}-$ 1 comparators) Pelgrom (2017).
However, the oscillatron design is intrinsically multifunctional in a temporal sense. When run for a time $t_{I}$, the outputs of the set of oscillators at all times less than $t_{I}$ provide the binary representation of all analog numbers less than $I$. Thus $n$ oscillators and $n$ comparators observed at $m$ different times would allow the analog-to-$n$-bit digital conversion of $m$ analog inputs, within a single run of the set of oscillators. The mean time required for the computation of each analog number then scales as $(2^{n}-1)/m$, while for a successive approximation register it is $mn$. For $m$ sufficiently large, the oscillator design is more efficient.
In this section, for simplicity, we set the damping coefficient $\gamma$ to zero. The presence of damping will cause the magnitude of the oscillator to decay over time, limiting the time for which the oscillator’s departure from zero can be measured, and so limiting the size of the analog number $I$ that can be represented. To counter this decay, energy can be input to the oscillator.
III Training oscillatrons by gradient descent to perform multifunctional computation